# Improving Complex Reasoning over Knowledge Graph with Logic-Aware Curriculum Tuning

### v1

@@ Abstract
Answering complex logical queries over incomplete knowledge graphs (KGs) is challenging. Most previous works have focused on learning entity/relation embeddings and simulating first-order logic operators with various neural networks. However, they are bottlenecked by the inability to share world knowledge to improve logical reasoning, thus resulting in suboptimal performance. In this paper, we propose a complex logical reasoning schema over knowledge graphs upon large language models (LLMs), containing a curriculum-based logical-aware instruction tuning framework, named LACT. Specifically, we augment the arbitrary first-order logical queries via binary tree decomposition, to stimulate the reasoning capability of LLMs. To address the difficulty gap among different types of complex queries, we design a simple and flexible logic-aware curriculum learning framework. Experiments across widely used datasets demonstrate that LACT has substantial improvements (brings an average +5.5% MRR score) over advanced methods, achieving the new state-of-the-art. Our code and model will be released at GitHub and huggingface soon.
1 Introduction
@@
Large-scale knowledge graphs (KGs) such as FreeBase  (?) stores structural knowledge in a collection of fact triplets and are widely adopted by many domains. Unfortunately, KGs are often incomplete, leaving many missing triplets undiscovered. Thus, complex logical reasoning over such KGs  (??) is challenging and has attracted much attention in the recent years. A complex logical query can be represented with First-Order Logic (FOL) that includes logical operators such as conjunction ()($\wedge$)( ∧ ), disjunction ()($\vee$)( ∨ ), negation (¬)($\neg$)( ¬ ), and existential quantifier ()($\exists$)( ∃ ), etc. A more direct approach involves the representation of computation graphs as Directed Acyclic Graphs (DAGs), which can be resolved through the systematic traversal of Knowledge Graphs (KG). This process entails the allocation of suitable entities to intermediate variables based on their structural attributes (?).
@@
Figure 1: Schematic illustration: a) Answering logical query over KG with LACT. b) The framework of Logic-Aware Curriculum Tuning over LLama. We leverage binary tree decomposition strategy (Seen in Section 4.2) to construct logic-rich FT corpus and Curriculum learning strategy (Seen in Section 4.3) to fine-tune a base LLM. c) Performing reasoning using well-designed prompts.
Inspired by the success of knowledge graph embedding (KGE) (??), a line of research proposes to answer complex logical queries by learning query embedding and simulating logical operators with well-designed neural networks  (????) Current research based on embeddings primarily focuses on the creation of diverse latent space geometries, such as vectors (?), boxes (?), hyperboloids (?), and probabilistic distributions (?), to effectively capture the semantic position and logical coverage of knowledge graph entities.
However, these approaches are limited in their performance due to the following. (1) Limited information: The information contained in a knowledge graph is usually incomplete and limited. When only the information from the knowledge graph can be used, it is difficult to answer some complex reasoning that lacks relevant information. (2) High complexity of logical queries: The intricacies of world knowledge determine the complexity of reasoning in practical applications, which determines that it is difficult to model the relationship of world knowledge through simple space geometries figures that may lose potentially complex relationship information (?), thus limiting the effect of complex logical reasoning. (3) Generalizability: KGE method for a particular KG can not generalize to other KGs which limits the applicability of these approaches in real-world scenarios where KGs can vary widely in terms of their structure and content.
@@
Recently, large language models (LLMs)  (???) show outperforming capabilities to a wide range of tasks (?????).  (??) construct prompt templates and apply LLMs as text-generators to answer complex queries. However, LLM without fine-tuning suffers from hallucination problem (?), especially when faced with such a knowledge-intensive task that generates answers on an incomplete KG rather than simple retrieval. Besides, previous tasks relied on manual classification of queries to improve performance, which is unrealistic in large-scale practical applications and also limits the types of queries. Finally, Previous methods generally decompose the problem into sub-problems, which greatly increases the cost of reasoning, especially considering that previous methods are generally based on closed-source models such as GPT. Therefore, compared to pure prompt engineering, we prefer to fine-tune our own model to solve the above problems.
@@
In this paper, we propose Logic-Aware Curriculum Tuning (LACT  [Uncaptioned image]), a novel fine-tune framework for answering complex logical query, which stimulates the ability of LLMs to perform complex reasoning on knowledge graphs. We propose a strategy to incorporate the knowledge contained in the KGs into our training corpus to activate the corresponding knowledge of the LLMs and supplement the missing relevant knowledge of the LLMs during the fine-tuning process. At the same time, we have proven that data argument by binary tree decomposition can arouse the corresponding capabilities of LLMs and effectively improve their reasoning performance. At last, we show that curriculum learning(?) can effectively smooth the difficulty differences between different types of queries and greatly improve the results of difficult queries. In summary, our contribution is three-folded:
@@
  • We propose a logic-aware curriculum fine-tuning (LACT) paradigm for complex logical reasoning over KGs.
@@
  • LACT achieves state-of-the-art performance beyond embedding-based and PLM-based methods, using only a 7B llama2 model.
@@
  • Through extensive experimentation, we found that fine-tuning corpus constructed with rigorous logical context over KGs and curriculum learning can significantly enhance LLM logical reasoning ability.
2 Related Works
2.1 Logical Reasoning over Knowledge Graph
@@
Knowledge graphs can be widely used to enhance natural language reasoning (??), understanding (??), and generation (??). Given a FOL query over a KG, complex logical reasoning aims to answer correct entities, which contains not only multi-hop paths but logical operators(??). Most of current approaches concentrated on learning meaningful query embeddings (?????). Neuralizing logical operators through a specific embedding space, thereby embedding FOL queries into a vector space(???), or probabilistic distribution(?), and predict answers by locating nearest neighbours to answer set representation. Additionally, approaches such as CQD(?) have focused on improving the performance of complex reasoning tasks through the answer composition of simple intermediate queries, and QTO (?) proposes query computation tree optimization that can efficiently find the exact optimal solutions. However, embedding-based methods usually lack interpretability as there is no explicit mapping between the embedding and the set of entities, and this limits their generalization ability to more complicated query structures.
@@
Moreover, PLM-based methods consider complex logical reasoning as text-based tasks by modeling pre-trained language models. The short textual descriptions are organized as an input sequence and encoded by the PLMs (???). However, limited by the performance limitations of the base model, the PLM method has always been deficient in understanding world knowledge and reasoning capabilities, limiting its performance in complex reasoning.
2.2 LLMs for KG Reasoning
@@
In recent years, substantial advancements have been witnessed in the domain of LLMs (????). Among these, instruction tuning (IT) (?) and the alignment of the model (?) with human preferences stand out.
@@
Within the realm of LLM, the integration of LLMs with Knowledge Graphs (KG) (???) constitutes a prominent and consequential research avenue.
@@
Leveraging its potent generative capabilities, LLMs prove invaluable in addressing Knowledge Graph-related tasks, including but not limited to Knowledge Graph Completion (KGC) (??), entity alignment (?), Knowledge Graph Question Answering (KGQA) (?), and others (?). Consequently, the synergy between Knowledge Graphs for LLMs (KG4LLM) and LLMs for Knowledge Graphs (LLM4KG) emerges as an essential focal point, bearing significance in advancing the collective capabilities of both entities.
@@
We focus on applying LLMs in the Complex Logical Reasoning task, which has not been carefully studied yet. (?) made the initial attempt by prompt engineer but it lacks in-depth research and simply uses LLMs as text generators.
2.3 Curriculum Learning
@@
The idea of training a neural network in an easy-to-difficult fashion can be traced back to  (??).Based on these works, Curriculum Learning is first proposed in (?). They designed several toy experiments to demonstrate the benefits of curriculum strategy both in image classification and language modeling. When focusing on the field of NLP, by experimenting with several heuristics,  (??) migrated the success of CL to NLU tasks. (??) improved the machine translation modeling by carefully designing different curriculum. Recently, with the rise of LLM, (?) discovered the huge potential of CL in in-context learning, while (?) focus on the improvement of CL for LLM in pre-train. However, we are committed to exploring the huge potential of CL in fine-tuning LLM.
3 Preliminary
3.1 Knowledge Graph
@@
In our work, a knowledge graph is $\mathcal{G}=(\mathcal{E,R,T})$ where $\mathcal{E,R}$ are the set of entity and relation respectively. With regard to generality, KG can be organized as a set of triplets $\{\mathcal{T}=(v_{s},r,v_{t})|v_{s},v_{t}\in\mathcal{E},r\in\mathcal{R}\}$, where $v_{s}/v_{t}$ denotes the head/tail entity.
3.2 Complex logical query
@@
Complex logical query is used for retrieving and manipulating data stored in knowledge graphs, which is grounded in a subset of FOL. The process of answering a complex logical query involves trying to match a suitable results using the composition of queries:
@@
where $q$ denotes a FOL query. Note that Eq. (1) is conjunctive normal form (CNF) and Eq. (2) is disjunctive normal form (DNF). The two can be equivalently converted to each other via De Morgan’s law. Following previous works (?), we focus on modeling the operations: projection $r(\cdot)$, conjunction ($\wedge$), disjunction ($\vee$), and negation ($\neg$). Additionally, note that existential positive first-order (EPFO) queries only includes projection, conjunction ($\wedge$), and disjunction ($\vee$).
4 Methodology
4.1 Instruction Tuning on LLMs
In this section, we introduce how to incorporate the KG information in the text-based prompt.
When applying LLMs to complex logical reasoning, we denote an LLM as $\mathcal{M}$ which is a text decoder to generate the corresponding output. If we start from the above definition, this task can be modelled as a text generation task. However, triplet generation is different from vanilla text generation because the entities and the relation in the triplet prompt have complex semantic information defined by the given KG. In fact, we want the generated answers to be entities that exist in KG itself. Without these knowledge, the predicted answers are unreliable and unstable. Thus, incorporating the KG information into the prompt to provide more auxiliary information is the key to engage LLMs in complex logical reasoning.
In particular, when we fine-tune $\mathcal{M}$, we can treat the training corpus as a set of question-answer pairs $(\mathcal{S},\mathcal{A})$. For the task of complex logical reasoning over knowledge graph, the input textual sequence $\mathcal{S}$ consists of the description of question $\mathcal{D}$, knowledge graph neighbourhood information(i.e. related triplets) $\mathcal{X}$ and logical query. In our work, we used a simple but effective method called greedy depth traversal algorithm to search for neighbourhood information and we simply discarded all samples that exceeded the token limit. (Detailed algorithm and token distribution can be found in Appendix A). The logical query contains the textual information about the query $q_{\tau}$ with the query structure $\tau$ and specific query content $Q_{\tau}$ that needs to be processed, which can be denoted as $f_{1}(q_{\tau})$. Likewise,we can denote the output answer A as $f_{2}(V_{\tau})$, where $f_{2}$ indicates textualization of $V_{\tau}$ here. In summary, the fine-tune training corpus $\mathcal{C}$ can be expressed in the following form:
@@ 4.1 Instruction Tuning on LLMs
The model $\mathcal{M}$(parameterized by $\theta$) is fine-tuned by the next token prediction task. We fine-tune $\mathcal{M}$ to obtain the final model by maximizing the log-likelihood of the next token. The training objective can be formulated as
@@
4.2 Data Augmentation by Binary Tree Decomposition
@@
Chain of thought (COT) enables models to decompose multi-step problems into intermediate steps, subsequently improving the reasoning abilities of LLMs (?). However, pure-prompt based reasoning needs more incontext memory to perform complex logical reasoning. Considering complex logical queries, which query structure can be transferred into the form of a DAG and its hierarchical structure becomes a natural fit for decomposition into a series of sub-problems. So we propose a method for data augment based on Binary Tree Decomposition Mechanism to stimulate LLMs with the potential to decompose a complex query into a chain of simple queries. Binary Tree Decomposition Mechanism. The Binary Tree Decomposition Mechanism is divided into the following three steps:
Query Computation Tree. For a complex FOL query, like the example shown in Figure 1, its computation graph that is a directed acyclic graph can be converted into a tree where the root node is $v_{?}$. query computation tree. The answer variable and the constant entities in the query correspond to root and leaf nodes in the query computation tree. Each edge in the query computation tree points from the child node to the parent node. It can be recursively deduced that the subtree rooted at any non-leaf node in the tree corresponds to a subquery. In Appendix B, we provide a systematical procedure for transforming an FOL query to its query computation tree.
Binary Tree Decomposition. For the one-to-many intersection/union structures in the tree, we separate each parent node into two child nodes. Note that the union branches merging step may create one-to-many structures that consist of both intersection and union edges, take Figure 1 for an example. This can be taken care of by first separating $v_{?}$ into an intersection structure ($v_{3}^{\prime}$ and $v_{5}^{\prime}$ in the example), and then separating the child node into an intersection structure ($v_{1}^{\prime}$ and $v_{2}^{\prime}$ for $v_{3}^{\prime}$, $v_{4}^{\prime},v_{3}$ for $v_{5}^{\prime}$), where $v^{\prime}$ denotes an intermediate entity retrieved by Neighborhood Retrieval Algorithm (Seen in Appendix A).
Reverse Level Traversal. Finally, we decompose the binary computation tree into independent branches. Since the root node of the calculation tree is the answer entity, we perform a hierarchical traversal of all non-leaf nodes of the binary tree in reverse. As shown in Figure 1, the complex FOL query is decomposed into a sequence:
@@
Data Augmentation. Now we can turn any loopless FOL query into a series of separate subqueries. We use a defined template to integrate the decomposition process into the answers to the training corpus. So, the training corpus $\mathcal{C}$ can be transferred into the following form:
@@
4.3 Fine-tuning Enhanced by Curriculum Learning
As mentioned in previous sections, though decomposing into chain responses, complex queries still vary greatly in difficulty and complexity due to differences between query structure.
Naturally, we believe that these different types of samples should not be simply lumped together. Intuitively, we incorporate curriculum learning into our training. To be specific, In view of the particularity of complex reasoning data, when we decompose it into logical chains, naturally, we can use the number of decomposed sub-logical queries as a natural difficulty discriminator to select different types of queries, e.g., a 1p query would be defined as difficulty-1, while a 2p query, which can be decomposed into two projection queries and an intersection query, would be defined as difficulty-3. The detailed difficulty discriminating process will be shown in Appendix (Table S1).

### v4

@@ Abstract
Answering complex queries over incomplete knowledge graphs (KGs) is a challenging task. Most previous works have focused on learning entity/relation embeddings and simulating first-order logic operators with various neural networks. However, they are bottlenecked by the inability to share world knowledge to improve logical reasoning, thus resulting in suboptimal performance. In this paper, we propose a complex reasoning schema over KG upon large language models (LLMs), containing a curriculum-based logical-aware instruction tuning framework, named LACT. Specifically, we augment the arbitrary first-order logical queries via binary tree decomposition, to stimulate the reasoning capability of LLMs. To address the difficulty gap among different types of complex queries, we design a simple and flexible logic-aware curriculum learning framework. Experiments across widely used datasets demonstrate that LACT has substantial improvements (brings an average +5.5% MRR score) over advanced methods, achieving the new state-of-the-art.
Codehttps://github.com/TianleXia0914/LACT
Introduction
@@
Industrial-scale knowledge graphs (KGs) like FreeBase store structural knowledge in a collection of fact triplets and are widely adopted by many domains. Unfortunately, KGs are often incomplete, leaving many missing triplets undiscovered. Thus, complex logical reasoning over such incomplete KGs (Hamilton et al. 2018; Bai et al. 2023) is a challenging task and has attracted much attention in the recent years. In the realm of complex logical queries, Existential First Order queries with a single free variable ($\bm{EFO_{1}}$) emerge as a powerful tool that includes logical operators such as the basic projection operation ()($\exists$)( ∃ ) and the corresponding negation operation (¬)($\neg$)( ¬ ), two relational operation intersection ()($\wedge$)( ∧ ) and union ()($\vee$)( ∨ ), etc. A very straightforward approach is to represent the computation graph as a Directed Acyclic Graph (DAG), which can be addressed by a systematic traversal of a knowledge graph (KG). This process entails the allocation of suitable entities to intermediate variables based on their structural attributes (Dalvi and Suciu 2007).
@@
Figure 1: Schematic illustration: a) Answering logical query over KG with LACT. b) The framework of Logic-Aware Curriculum Tuning over LLama. We leverage the binary tree decomposition strategy (Seen in Section Methodology) to construct a logic-rich FT corpus and the Curriculum learning strategy (Seen in Section Methodology) to fine-tune a base LLM. c) Performing reasoning using well-designed prompts.
In the wake of knowledge graph embedding (KGE)’s resounding success, (Bordes et al. 2013; Bai et al. 2021), a series of work seeks to answer complex logical queries by learning query embedding and simulating logical operators with well-designed neural networks  (Chen, Hu, and Sun 2022; Zhu et al. 2022; Zhang et al. 2021; Arakelyan et al. 2020). In contemporary research, the exploration of KGE method predominantly centers around the intricate design of various embedding representations, including geometric embedding like Euclidean geometry modeling (Hamilton et al. 2018; Ren, Hu, and Leskovec 2019) and Riemannian geometry modeling (Choudhary et al. 2021), and embedding based on probability modeling (Ren, Hu, and Leskovec 2019), seeking to encapsulate the semantic representation and vector mapping of KGs’ entities and relations.
However, the embedding-based approach described above also has some limitations. (1) Limited information: Knowledge graphs often contain incomplete and insufficient information. When only the information from the knowledge graph is available, it becomes challenging to perform complex reasoning that requires additional or missing information. (2) Insufficient utilization of semantic information: Current knowledge graph embedding (KGE) methods primarily focus on the structural information of the knowledge graph, paying little attention to the actual semantics of entities and relationships. (3) High complexity of logical queries: he intricate nature of real-world knowledge contributes to the complexity of reasoning in practical applications. As a result, it is challenging to accurately model the relationships within real-world knowledge using simple geometric representations. Such simplifications may overlook potentially complex relational information (Choudhary et al. 2021), thereby limiting the effectiveness of advanced logical reasoning. (4) Generalizability: A Knowledge Graph Embedding (KGE) model tailored to a specific knowledge graph cannot generalize across different knowledge graphs. This limitation hinders the practical application of such approaches in real-world scenarios, where knowledge graphs often vary significantly in both structure and content.
@@
Recently, large language models (LLMs)  (Achiam et al. 2023; Touvron et al. 2023a) have showed outperforming capabilities for a wide range of tasks (Ouyang et al. 2022; Zhong et al. 2023b; Peng et al. 2023; Lu et al. 2023; Guo et al. 2025). With the rise of this trend, (Choudhary and Reddy 2023; Liu et al. 2024a) construct prompt templates and apply LLMs as text-generators to answer complex queries. However, LLM without fine-tuning suffers from hallucination problem (Zhang et al. 2023c), especially when faced with such a knowledge-intensive task that generates answers on an incomplete KG rather than simple retrieval. Besides, previous tasks relied on manual classification of queries to improve performance, which is unrealistic in large-scale practical applications and also limits the types of queries. Finally, Previous methods generally decompose the problem into sub-problems, which greatly increases the cost of reasoning, especially considering that previous methods are generally based on closed-source models such as GPT. Therefore, compared to pure prompt engineering, we prefer to fine-tune our model to solve the above problems.
@@
In this paper, we propose Logic-Aware Curriculum Tuning (LACT[Uncaptioned image]), a novel fine-tune framework for answering complex logical query, which stimulates the ability of LLMs to perform complex reasoning on knowledge graphs. We propose a strategy to incorporate the knowledge contained in the KGs into our training corpus to activate the corresponding knowledge of the LLMs and supplement the missing relevant knowledge of the LLMs during the fine-tuning process. At the same time, we have proven that data argument by binary tree decomposition can arouse the corresponding capabilities of LLMs and effectively improve their reasoning performance. At last, we show that curriculum learning (Bengio et al. 2009) can effectively smooth the difficulty differences between different types of queries and greatly improve the results of difficult queries. In summary, our contribution manifests in three distinct facets:
@@
  • We propose a Logic-Aware Curriculum tuning (LACT) paradigm for complex logical reasoning over KGs.
@@
  • LACT achieves state-of-the-art performance beyond embedding-based and PLM-based methods, using only 7B models.
@@
  • Through extensive experimentation, we found that SFT Corpus constructed by queries on incomplete KGs via Binary Tree Decomposition and Curriculum Learning can significantly enhance LLM logical reasoning ability.
Related Works
Logical Reasoning over Knowledge Graph
@@
Knowledge graphs can be widely used to enhance natural language reasoning (Zhang and Yao 2022; Zhao et al. 2023), understanding (Zhong et al. 2023a; Liu et al. 2023), and generation (Ding, Wang, and Liu 2023; Pan et al. 2024b). Given a $EFO_{1}$ query over a KG, complex logical reasoning aims to answer the correct entities, which contains both multi-hop queries and logical operators(Guu, Miller, and Liang 2015; Hamilton et al. 2018) over a incomplete KG. Most of the current approaches have focused on learning meaningful KG embeddings including (Arakelyan et al. 2020; Chen, Hu, and Sun 2022; Zhang et al. 2021; Zhu et al. 2022; Wang et al. 2023b). Neuralizing logical operators through a specific embedding space, thereby embedding $EFO_{1}$ queries into a multi-dimensional vector space(Hamilton et al. 2018; Ren, Hu, and Leskovec 2019), or into a specific probability distribution(Ren, Hu, and Leskovec 2019; Choudhary et al. 2021), then obtain the final answer set by fitting the nearest neighbour representation based on relational operations. Additionally, recent work like CQD (Arakelyan et al. 2020) its optimization work (Arakelyan et al. 2024) improves performance by working to reduce query difficulty by decomposing complex queries into simple one-hop queries, QTO (Bai et al. 2023) introduces query computation tree optimization and LMPNN (Wang et al. 2023f) aggregates logical messages passed on the query graph by training an MLP network. Some work (Yin, Wang, and Song 2023) has begun to expand into the area of cyclic query. Despite their efficacy, embedding-based methods often suffer from a lack of generalization due to their specificity in embedding knowledge graphs. Meanwhile, this limitation on the embedding range hinders their ability to effectively generalize to more complex query structures.
@@
Moreover, PLM-based methods view complex logical reasoning as text-generation tasks by modeling pre-trained language models. $EFO_{1}$ queries are organized into input-output sequence pairs after textualization and encoded by PLM (Wang et al. 2023c; Xu et al. 2023; Wang et al. 2023c). However, limited by the performance limitations of the base model, the PLM method has always been deficient in understanding world knowledge and reasoning capabilities, limiting its performance in complex reasoning.
LLMs for KG Reasoning
@@
In recent years, substantial advancements have been witnessed in the domain of LLMs (Achiam et al. 2023; Touvron et al. 2023b; Peng et al. 2023; Zhong et al. 2023b; Zeng et al. 2022). Among these, instruction tuning (IT) (Ouyang et al. 2022) and the alignment of the model (Wang et al. 2023e) with human preferences stand out.
@@
Within the realm of LLM, the integration of LLMs with Knowledge Graphs (KG) (Pan et al. 2024a; Wang et al. 2023a; Luo et al. 2024) constitutes a prominent and consequential research avenue.
@@
Leveraging its potent generative capabilities, LLMs prove invaluable in addressing Knowledge Graph-related tasks, including but not limited to Knowledge Graph Completion (KGC) (Zhu et al. 2023; Zhang et al. 2023b), entity alignment (Zhang et al. 2023a), Knowledge Graph Question Answering (KGQA) (Luo et al. 2024), and others (Luo et al. 2023). Consequently, the synergy between Knowledge Graphs for LLMs (KG4LLM) and LLMs for Knowledge Graphs (LLM4KG) emerges as an essential focal point, bearing significance in advancing the collective capabilities of both entities.
@@
Our research endeavours center around the utilization of Language Models (LLMs) for the Complex Logical Reasoning task, an area that remains relatively unexplored. (Choudhary and Reddy 2023) made the initial attempt by prompt engineer but it lacks in-depth research and simply uses LLMs as text generators.
Curriculum Learning
@@
The concept of progressively training neural networks from simple to complex configurations has its origins dating back to (Elman 1993; Krueger and Dayan 2009). Based on these works, Curriculum Learning is first proposed in (Bengio et al. 2009). A series of illustrative experiments were meticulously crafted to showcase the advantages of employing curriculum strategies in both image classification and language modeling. When focusing on the field of NLP, by experimenting with several heuristics,  (Sachan and Xing 2016; Xu et al. 2020) migrated the success of CL to NLU tasks. (Ding et al. 2021; Zhou et al. 2021) improved the machine translation modeling by carefully designing different curricula. Recently, with the rise of LLM, (Liu et al. 2024b) discovered the huge potential of CL in in-context learning, while (Wang et al. 2023d) focus on the improvement of CL for LLM in pre-train. However, we are committed to exploring the huge potential of CL in fine-tuning LLM.
Preliminary
Knowledge Graph
@@
In our work, a knowledge graph is $\mathcal{G}=(\mathcal{E,R,T})$ where $\mathcal{E,R}$ are the set of entity and relation respectively. With regard to generality, KG can be considered as a collection of triplets $\{\mathcal{T}=(v_{s},r,v_{t})|v_{s},v_{t}\in\mathcal{E},r\in\mathcal{R}\}$, where $v_{s}/v_{t}$ denotes the head/tail entity.
Complex logical query
@@
Complex logical query is used for retrieving and manipulating data stored in knowledge graphs, which is grounded in a subset of $EFO_{1}$ query. The process of answering a complex logical query involves trying to match suitable results using the composition of queries:
@@
where $q$ denotes a $EFO_{1}$ query and $q_{i}$ denotes a sub-query. Note that Eq. (1) is a conjunctive normal form (CNF) and Eq. (2) is a disjunctive normal form (DNF) (Ren, Hu, and Leskovec 2019). The two can be equivalently converted to each other via De Morgan’s law.
Methodology
Instruction Tuning on LLMs
Within this section, we elucidate the methodology for seamlessly integrating Knowledge Graph (KG) information into the textual prompt.
In the context of complex logical reasoning, we designate a Language Model (LLM) as $\mathcal{M}$ serving as a text decoder to produce the corresponding output. By commencing with the aforementioned definition, we can frame this task as a text generation problem. n contrast to vanilla text generation, triplet generation involves more complexity due to the intricate semantic information associated with the entities and relations in the triplet prompt, as defined by the given knowledge graph (KG). In fact, we want the generated answers to be entities that exist in KG itself. Without this knowledge, the predicted answers are unreliable and unstable. Thus, to engage LLM in complex logical reasoning, a crucial step involves incorporating knowledge graph (KG) information into the prompt. This integration provides additional auxiliary context, effectively enhancing the performance of LLM.
In particular, when we fine-tune $\mathcal{M}$, we can treat the training corpus as a set of question-answer pairs $(\mathcal{S},\mathcal{A})$. When considering complex reasoning over KGs, the input textual sequence $\mathcal{S}$ consists of the description of question $\mathcal{D}$, knowledge graph neighbourhood information(i.e. related triplets) $\mathcal{X}$ and logical query. In our work, we used a simple but effective method called greedy depth traversal algorithm to search for neighbourhood information and we simply discarded all samples that exceeded the token limit. (Detailed algorithm and token distribution can be found in Appendix). The logical query contains the textual information about the query $q_{\tau}$ with the query structure $\tau$ and specific query content $Q_{\tau}$ that needs to be processed, which can be denoted as $f_{1}(q_{\tau})$. Likewise,we can denote the output answer A as $f_{2}(V_{\tau})$, where $f_{2}$ indicates textualization of $V_{\tau}$ here. In summary, the fine-tune training corpus $\mathcal{C}$ can be expressed in the following form:
@@
The model $\mathcal{M}$(parameterized by $\theta$) is fine-tuned by the next token prediction task. We fine-tune $\mathcal{M}$ to obtain the final model by maximizing the log-likelihood of the next token. The training loss can be formulated as
@@
Data Augmentation by Binary Tree Decomposition
@@
Chain of thought (COT) enables models to decompose multi-step problems into intermediate steps, subsequently improving the reasoning abilities of LLMs (Wei et al. 2022). However, pure-prompt-based reasoning needs more in-context memory to perform complex logical reasoning. Considering complex logical queries, which query structure can be transferred into the form of a DAG and its hierarchical structure becomes a natural fit for decomposition into a series of sub-problems. So we propose a method for data augment based on Binary Tree Decomposition Mechanism to stimulate LLMs with the potential to decompose a complex query into a chain of simple queries. Binary Tree Decomposition Mechanism. The Binary Tree Decomposition Mechanism is divided into the following three steps:
Query Computation Tree. Existing embedding-based methods rely on DNF to avoid embedding divergence (Ren and Leskovec 2020). Our method bypasses logical operator learning. Consequently, this allows us to convert any query into a logic-equivalent tree form through iterative sub-query recursion. For a complex $EFO_{1}$ query, like the example shown in Figure 1, its computation graph that is a directed acyclic graph can be converted into a tree where the root node is $v_{?}$. The tail entities and intermediate entities in complex queries are in one-to-one correspondence with the root nodes and leaf nodes in the generated calculation tree respectively. In view of the reverse order that the tail entity brings to the root node, the edges in the tree corresponding to the relationship in each query are child-to-parent directions. In the Appendix, We demonstrate a simple but systematic method for converting complex queries into computational trees.
Binary Tree Decomposition. We split each 1-to-n intersection/union node into n corresponding child nodes. Consider that merging union branches may result in a 1-to-n structure consisting of intersection and union edges, as shown in Figure 1. This can be properly addressed by separating $v_{?}$ into an intersection node structure ($v_{3}^{\prime}$ and $v_{5}^{\prime}$ in Figure 1) firstly, then decomposing the node that was decomposed in the previous layer into a deeper intersection node structure ($v_{1}^{\prime}$ and $v_{2}^{\prime}$ for $v_{3}^{\prime}$, $v_{4}^{\prime},v_{3}$ for $v_{5}^{\prime}$), where $v^{\prime}$ denotes an intermediate entity retrieved by Neighborhood Retrieval Algorithm (Seen in Appendix).
Reverse Level Traversal. Finally, we decompose the binary computation tree into independent branches. Since the root node of the calculation tree is the answer entity, we perform a hierarchical traversal of all non-leaf nodes of the binary tree in reverse. As shown in Figure 1, the complex $EFO_{1}$ query is decomposed into a sequence:
@@
Data Augmentation. Now we can turn any loopless $EFO_{1}$ query into a series of separate subqueries. We use a defined template to integrate the decomposition process into the answers to the training corpus. So, the training corpus $\mathcal{C}$ can be transferred into the following form:
@@
Fine-tuning Enhanced by Curriculum Learning
As mentioned in previous sections, though decomposing into chain responses, complex queries still vary greatly in difficulty and complexity due to differences between query structures.
Naturally, we believe that these different types of samples should not be simply lumped together. Intuitively, we incorporate curriculum learning into our training. To be specific, given the particularity of complex reasoning data, when we decompose it into logical chains, naturally, we can use the number of decomposed sub-logical queries as a natural difficulty discriminator to select different types of queries, e.g., a 1p query would be defined as difficulty-1, while a 2p query, which can be decomposed into two projection queries and an intersection query, would be defined as difficulty-3. The detailed difficulty discriminating process will be shown in the Appendix.